OpenStudy (anonymous):

May you help me with this problem? @ganeshie8 @phi

OpenStudy (anonymous):

Look at the ratios of each side - what did you get...for example 22/33 = 2/3 you solve the other two sides

jhonyy9 (jhonyy9):

so think the first one because 22/14 = 33/21 or 14/24 = 21/36 or 24/22 = 36/33 hope these will help you understanding sure easy right